OpenStudy (anonymous):

how do i express each equation in slope-intercept form? x + 6/12y = -83 12x + 4y = -1772

OpenStudy (radar):

Arrange the equation to the y=mx+b form 1/2 y=-x-83 y=-2x-166 for the first one.

OpenStudy (radar):

4y=-12x-1772 y=-3x-443 for the other.

OpenStudy (anonymous):

can you please explain the process? thanks <3

OpenStudy (radar):

You the rules of algebra. You want to get y on the left and the x + b on the right. Now the first equation: x+ (6/12)y=-83 . Simplify and use the rules of algebra. First simplify the 6/12 which is 1/2. Next subtract x from both sides you now have x-x+(1/2)y=-x-83 simplify: 1/2 (y)=-x-83 Now solve for y, by multiplying boths sides by 2 getting: y=-2x-166. You now have it in the desired form The slope (m) is -2 and is the coefficient of x. constant b is -166 and is the y intercept (the value of y when x is 0)

OpenStudy (radar):

Any further questions chaplaps?

OpenStudy (anonymous):

no, not as of yet. :) thanks

OpenStudy (radar):

Good luck.

OpenStudy (anonymous):

the second equation is confusing, how can I solve it ? im stuck at 4y = -12x-1772